# The smallest open polynomial Diophantine equations

### v2

@@ 3 Describing all integer/rational solutions
with $11/p+1/q+1/r<$ 1. An integer solution $(x,y,z)$ to (18) is called primitive if $\text{gcd}(x,y,z)=1$, and it is known that under condition $11/p+1/q+1/r<$ 1 the set of primitive solutions is finite. Table 4 presents the smallest equations of the form (18) for which the author of this document does not know the complete list of their primitive solutions.
@@
$40$ $x^{4}+2y^{3}+z^{3}=0$ $56$ $x^{4}+3y^{3}+2z^{3}=0$ $56$ $x^{5}+2y^{3}+z^{3}=0$
$48$ $x^{4}+2y^{3}+2z^{3}=0$ $56$ $x^{4}+4y^{3}+z^{3}=0$ $56$ $x^{5}+y^{4}-2z^{2}=0$
$48$ $x^{4}+3y^{3}+z^{3}=0$ $56$ $02x^{4}+2y^{3}+z^{3}=$ 0 $60$ $x^{5}+y^{4}-3z^{2}=0$
$48$ $02x^{4}+y^{3}+z^{3}=$ 0 $56$ $x^{4}-y^{4}+3z^{3}=0$ $60$ $x^{5}+y^{4}+3z^{2}=0$
$48$ $x^{4}-y^{4}+2z^{3}=0$ $56$ $02x^{4}-y^{4}+z^{3}=$ 0
Table 4: Open equations of the form (18) with $11/p+1/q+1/r<$ 1 of size $H\leq 60$.
@@ 4 Existence of arbitrary large integer solutions
Given a homogeneous polynomial $P(x_{1},\dots,x_{n})$ with integer coefficients, determine whether equation (1) has an integer solution such that all variables are different from 00.
@@
$04x^{5}+4y^{5}+11z^{5}=$ 0 (21)
of size $H=608$. The next smallest is the equation
$7x^{4}-7y^{4}=25z^{4}$ (22)
of size $H=624$.
@@ 7 The shortest open equations
All equations in Table 15 have degree at least $4$. The shortest cubic equation for which Problem 6 is open is the equation
$17x^{3}+2y^{3}=3z^{2}+$ 1 (40)
@@ 8.1 Changes between version 1 and version 2
Equations $x^{4}+y^{4}+2z^{3}=0$ and $x^{5}+y^{4}+2z^{2}=0$ have been solved in the literature (see our forthcoming survey on generalized Fermat equations for details and references), hence these equations have been excluded from Table 4.
For the equations $02x^{2}-xyz-y^{2}-1=$ 0, $02x^{2}-xyz+y^{2}+1=$ 0 and $y^{2}+xyz+x^{3}+2=0$ Problem 2 has been solved111see https://mathoverflow.net/questions/449559/ and https://mathoverflow.net/questions/411958/. These equations have therefore been removed from Tables 5 and 13.
@@
For the equations $06x^{2}z+z^{2}y+3y^{3}+1=$ 0, $y^{2}+10xyz+x^{3}-x-2=0$ and $y^{2}+7xyz+3x^{3}-2=0$ Problem 6 has been solved333see https://mathoverflow.net/questions/466803/. Previously, these equations were the shortest cubic equations for which Problem 6 was open. These equations have now been removed from Section 7.

### v4

# A systematic approach to Diophantine equations: open problems

@@ 3 Describing all integer/rational solutions
with $1/p+1/q+1/r<1$. An integer solution $(x,y,z)$ to (18) is called primitive if $\text{gcd}(x,y,z)=1$, and it is known that under condition $1/p+1/q+1/r<1$ the set of primitive solutions is finite. Table 4 presents the smallest equations of the form (18) for which the author of this document does not know the complete list of their primitive solutions.
@@
$40$ $x^{4}+2y^{3}+z^{3}=0$ $56$ $x^{4}+4y^{3}+z^{3}=0$ $56$ $x^{5}+y^{4}-2z^{2}=0$
$48$ $x^{4}+3y^{3}+z^{3}=0$ $56$ $2x^{4}-y^{4}+z^{3}=0$ $60$ $x^{5}+y^{4}-3z^{2}=0$
$56$ $x^{4}+3y^{3}+2z^{3}=0$ $56$ $x^{5}+2y^{3}+z^{3}=0$ $60$ $x^{5}+y^{4}+3z^{2}=0$
Table 4: Open equations of the form (18) with $1/p+1/q+1/r<1$ of size $H\leq 60$.
@@ 4 Existence of arbitrary large integer solutions
Given a homogeneous polynomial $P(x_{1},\dots,x_{n})$ with integer coefficients, determine whether equation (1) has an integer solution such that all variables are different from 0.
@@
$4x^{5}+4y^{5}+11z^{5}=0$ (21)
of size $H=608$.
@@ 99 $z^{2}+y^{2}z+x^{3}y+1=0$ 99 $x^{3}y^{2}=z^{3}+2$ 99 $x^{4}+xy+y^{3}+1=0$
All equations in Table 15 have degree at least 44. The shortest cubic equation for which Problem 6 is open is the equation
$7x^{3}+2y^{3}=3z^{2}+1$ (39)
@@ 8.1 Changes between version 1 and version 2
Equations $x^{4}+y^{4}+2z^{3}=0$ and $x^{5}+y^{4}+2z^{2}=0$ have been solved in the literature (see [2] for details and references), hence these equations have been excluded from Table 4.
For the equations $^{2}-xyz-y^{2}-1=0, 2x2x$$^{2}-xyz+y^{2}+1=0$ and $y^{2}+xyz+x^{3}+2=0$ Problem 2 has been solved111see https://mathoverflow.net/questions/449559/ and https://mathoverflow.net/questions/411958/. These equations have therefore been removed from Tables 5 and 13.
@@
For the equations $6x^{2}z+z^{2}y+3y^{3}+1=0$, $y^{2}+10xyz+x^{3}-x-2=0$ and $y^{2}+7xyz+3x^{3}-2=0$ Problem 6 has been solved333see https://mathoverflow.net/questions/466803/. Previously, these equations were the shortest cubic equations for which Problem 6 was open. These equations have now been removed from Section 7.
8.2 Changes between version 2 and version 3
Equations $2x^{4}+y^{3}+z^{3}=0$, $x^{4}+2y^{3}+2z^{3}=0$, $x^{4}-y^{4}+2z^{3}=0$ and $x^{4}-y^{4}+3z^{3}=0$ have been solved (see [2] for details and references), hence these equations have been excluded from Table 4. Further, equation $2x^{4}+2y^{3}+z^{3}=0$ has been removed from Table 4, because it is equivalent to equation $x^{4}+4y^{3}+z^{3}=0$ presented in the same table.
8.3 Changes between version 3 and version 4
Equation $x^{2}y^{2}+x^{2}z-z^{2}-1=0$ reduces the equation to $u^{2}-4uxy-x^{4}+4=0$ after substitution $z\to xy+(x^{2}-u)/2$ for a new variable $u$. This equation is similar to ones removed between versions 1 and 2, and can be solved by the same method. Equation $y^{2}+z^{2}=x^{5}-1$ has been solved by Jeremy Rouse using a very simple argument, which is also applicable to $y^{2}+2x^{3}y+z^{2}+1=0$. All these equations have been deleted from Table 14.
Equation $7x^{4}-7y^{4}=25z^{4}$ has been listed as the next-smallest open of the form (20) after (21), but has now been solved by Nguyen Xuan Tho.